OpenStudy (anonymous):

A box is has dimensions of 14 inches long, 1.5 feet wide, and 7 inches high. What is the volume of the box? The formula for the volume is V = l • w • h. 1.02 cubic inches (in3) 12.25 cubic inches (in3) 147 cubic inches (in3) 1764 cubic inches (in3)

OpenStudy (anonymous):

If you want you can do it all in one step: V=l×w×h=14"×(1.5ft×12"1ft)×7"

OpenStudy (anonymous):

the desired answer is in cubic inches we convert 1.5 feet to inches: 1.5 feet * 12 inches/1 ft = 18 inches 14 * 18 * 7 = 1764 cubic inches